the weekly wages of John and Stephanie are in the ratio 5:3. their weekly expenditures are in the ration 9:5. each saves $30 dollars a week, find their weekly wages.
Answer by ankor@dixie-net.com(22740)

the weekly wages of John and Stephanie are in the ratio 5:3.
their weekly expenditures are in the ration 9:5. each saves $30 dollars a week, find their weekly wages.
:
Let x = the multiplier for their wages, then:
5x = J's wages
3x = S's wages
;
Let y = the multiplier for their expenditures, then
9y = J's expend
5y = S's expend
:
Write an equation for each
5x - 9y = 30
3x - 5y = 30
:
Use elimination here multiply the 1st equation by 3, the 2nd by 5
15x - 27y = 90
15x - 25y = 150
------------------subtraction eliminates x, find y
-2y = -60
y = +30 is the multiplier for expenses
Find x
5x - 9(30) = 30
5x - 270 = 30
5x = 30 + 270
5x = 300
x = 300/5
x = 60 is the multiplier for wages
Then
5(60) = $300 is J's wages
3(60) = $180 is S's wages